on a right triangle side A has a lenghth of 6 inches.Angle Y is 45 degrees. What is the length of the hypotenuse. I know the other missing angle is 45 degrees. I used the sin function on my calculater with 45 degrees and came up with 17.97 inches. Im not shure if this is right or weather i should use the second power function as well. Please help.
Answer by stanbon(75887)

on a right triangle side A has a length of 6 inches.
Angle Y is 45 degrees.
What is the length of the hypotenuse. I know the other missing angle is 45 degrees.
-------
I used the sin function on my calculater with 45 degrees and came up with 17.97 inches. Im not sure if this is right or wheather i should use the second power function as well.
---------------------
Draw the picture:
You have a right isosceles triangle.
Each leg is 6 inches.
---
Using Pythagoras to solve for "h", the hypotenuse.
h^2 = 6^2 + 6^2
h^2 = 2*6^2
h = 6*sqrt(2) = 8.485.. inches
=====================================
Cheers,
Stan H.